### v3

@@ Abstract
Foundation models applied to bio-molecular space hold promise to accelerate drug discovery. Molecular representation is key to building such models. Previous works have typically focused on a single representation or view of the molecules. Here, we develop a multi-view foundation model approach, that integrates molecular views of graph, image and text. Single-view foundation models are each pre-trained on a dataset of up to 200M molecules and then aggregated into combined representations. Our multi-view model is validated on 22 tasks, encompassing ligand-protein binding, molecular solubility, ADME and toxicity. We show that the multi-view models perform robustly and are able to balance the strengths and weaknesses of specific views. We then apply this model to screen compounds against over$>100$ G Protein-Coupled receptors (GPCRs). We prioritize 33 targets that are related to Alzheimer’s disease and employ our model to identify strong binders. They are validated through structure-based modeling and identification of key binding motifs.
@@ 1 Introduction
While sequence-based representations have been successful in capturing human-generated knowledge/information in LLMs, when it comes to representing natural objects like molecules, it is likely that sequence molecular representations alone will not be sufficient. Here we propose Multi-view Molecular Embedding with Late Fusion (MMELON) a flexible approach to aggregate multiple views in a foundation model setting and show that the resulting richer representations lead to better versatility across a large range of downstream tasks. The molecule use case demonstrates the effectiveness of the approach, which can be easily extended to include proteins and other biological entities. Our multi-view model is available on GitHub (https://github.com/BiomedSciAI/biomed-multi-view) and Hugging Face (https://Huggingface.co/ibm/biomed.sm.mv-te-84m).
@@
Graph representations are naturally conducive to capturing molecular symmetries and other structural properties in a flexible manner [18, 19, 20, 21, 2] However graph neural network (GNN) architectures lack the capacity to scale to large numbers of parameters and learn from very large datasets. Architectures that combine elements of graphs and transformer based approaches are better candidates for foundation models. Graph-transformers can represent graphs in a high-capacity neural network and several variants of this approach have already been explored [22, 23].
@@
Our multi-view molecular foundation model architecture, MMELON (Fig. 1A), leverages the suite of Biomedical Foundation Model (BMFM) technologies [30]. This multi-view architecture combines molecular representations to learn a joint base multi-modal embedding. This base embedding is then used to fine tune downstream tasks for chemical and biological property prediction. The model currently integrates three views: image, 2-dimensional chemically-bonded graph and text, although our architecture is extensible and can easily include other modalities such as 3-D conformations. For the image view we adopt ImageMol [25], a state of the art molecular foundation model trained on images of molecules. Graph and text views are each pre-trained on a 200M molecular dataset curated from public sources [31, 32]. Using a pre-trained aggregator module to combine the embeddings, we show that the individual modalities contain complimentary information. The weight of each view is a learned parameter, yielding a transparent interpretation. Our multi-view foundation model is fine-tuned on an extensive set of downstream tasks for molecular property prediction [33, 34, 35].
@@ 2.1 Overview of models and architectures
There are a number of multi-modal fusion strategies that may be employed to combine representations [47, 26, 27, 29]. We adopt a late fusion approach because it readily integrates separately pre-trained models and naturally supports analyses to gain insights into the role and importance of different modalities for different downstream tasks. In this approach, the singe-views are first separately trained and then combined using an aggregator sub-network (see Fig 1A). The single-view foundation models serve as pre-trained encoders that are utilized for both aggregator pre-training (see Sec. 4.5) and downstream task fine-tuning (see Section 2.3). For our aggregator architecture, we use an attention-based approach inspired by Ref. [28] where the contribution of the $m^{\text{th}}$ view to the overall embedding vector, $z_{i}^{\text{mv}}$, is weighted by coefficient $\alpha^{m}$,
@@ 2.2 Pre-trained models learn comprehensive embeddings
Image, Graph and Text offer intrinsically different views of chemical structure. However, it is expected that there is at least some overlap in their description of the embedding space. In order to characterize the pre-trained embeddings, we compute the correlation between their Euclidean distances in Fig. 1C. The embeddings are derived from a subset of $10,0000$ molecules sampled from the pre-training set. It can be seen that of the three single-view embeddings, Image is the most distinct. Text and Graph are highly correlated to each other ($c=0.7$), indicating that these two representations have a similar description of molecular similarity.
@@
Pre-trained molecular embeddings should ideally capture important aspects of the chemical structures of the underlying molecules. While there is no absolute way of measuring this, comparison with fingerprints that have been developed by domain experts can serve as a sanity check on whether pre-trained embeddings are in line with chemical knowledge  [11]. In Fig. 1D the correlation coefficients of the single-view models with four common fingerprint descriptors (Morgan, atom pair, MACCS and topological torsion) are plotted. [48, 13, 14, 15]. As a point of reference, we first compute the correlation between these fingerprints, showing that they are moderately correlated and yield a different description of the molecules (see Fig 1C). The atom pair fingerprint exhibits the highest correlation to the others.
@@
The correlations plotted in Fig. 1D are computed between the Euclidean distance of the pre-trained models and Tanimoto distance of the fingerprints over the same molecules used above. It can be seen that there is moderate correlation between pre-trained embedding and the fingerprints, demonstrating the validity of the embeddings. Overall, Text exhibits the highest correlation, followed by Graph and Image. Image is most correlated to the MACCS fingerprint, which is expected given that it is used in one of the ImageMol model’s pre-training tasks [25]. Finally, we note that the correlation with fingerprints only serves as a sanity check and strong agreement with fingerprints is not necessarily desirable for all down stream property prediction tasks, because each fingerprint is tailored using a limited set of hand-engineered features. A key objective of learning pre-trained multi-view embeddings is to achieve richer, more flexible representations.
@@
Figure 1: (A) Schematic of multi-view architecture. Embeddings from three single view pre-trained encoders (Image, Graph and Text) are combined by the aggregator module into a combined embedding. The network is fine-tunable for downstream predictive tasks. (B) Pre-training strategies for single view (Image, Text (sequence) and Graph) embeddings. These strategies are described in detail in Sec. 4.4. (C) The elements of the correlation matrix between the three-single view embeddings (lower triangle) and between four fingerprints (upper triangle). (D) The correlation between four fingerprints and the pre-trained, single-view embeddings.
@@ 2.3 Baseline performance: fine-tuning on downstream tasks
The prediction of electronic properties, ligand-protein binding affinity, pharmokinetic properties, and toxicity are examples of tasks that are essential to chemistry and drug discovery. MoleculeNet [34] is a collection of datasets covering such tasks that serves as a common test for machine learning models. Performance on this suite of benchmarks is neither decisive, nor necessarily reflective of all important use-cases in computational drug discovery, but currently it stands as the most widely used point of comparison between models. We also assess our model on predicting the inhibition of cytochrome P-450 isoforms [33](CYP) which are critical in drug metabolism pathways, and the recently released ComputationalADME dataset, which include tasks such as liver microsomal stability and plasma protein binding [49]. In addition to the choice of dataset, the choice of data splitting is critical in robust assessment of model performance. We primarily adopt scaffold-based methodologies, but also utilize random splits where appropriate (see SI for further discussion).
@@
We find that overall the multi-view model performs robustly across all tasks, spanning areas such as molecule solubility prediction (ESOL, Lipophilicity), binding activity assay-based classification (HIV, MUV) and ADME-Tox based tasks such as Tox21 and the ComputationalADME set (e.g. RPPB, HPPB, RLM, HLM). In particular, we test on cytochrome-P450 (CYP) inhibition for five CYP isoforms [33]. Performance as measured by ROC-AUC is high, ranging from $0.90$ to $0.82$ and results are comparable with SOTA performance [25]. For example, isoform CYP2C9 is one of most important contributor to drug metabolism [51] and our model achieved a high ROC-AUC value (0.90, Fig. 2A). The binding modes of top bio-active molecules computed from molecular docking align well with reported structures of CYP2C9 in complex with the drug losartan [52] (PDB ID: 8VX0, see Fig. 2B). In addition, the salient molecular moieties as predicted by the graph and image views of multi-view are both close to the catalytic site of CYP2C9 [53], suggesting our multi-view models could highlight the important binding features.
@@
One advantage of the multi-view late-fusion model architecture is that the contributions of single views to each task can be readily recovered through the aggregator weights, $\alpha$ (see Eqn. 3). The $\alpha$ values are visualized as a heat map for MoleculeNet tasks in results in Fig. 2C. We note that the pre-trained aggregator (see Sec. 4.5) is free to evolve during fine-tuning and adapt to the downstream tasks. The Graph model is heavily weighted for most tasks, which is consistent with the fact that Graph is overall the best performing single-view model. A complementary analysis measuring the agreement between different views of selected fine-tuned tasks is given in Sec. 4.7 and Figs. S4-S5. For CYP models, both Graph and Image are nearly equally correlated with the multi-view model, and this degeneracy can be related to trends in Fig. 2C.
With the exception of QM7, the Text Model is a minor contributor to the multi-view fine-tuned embeddings and while the predictions between text and multi-view are correlated, they exhibits the weakest relation (see Figs. S4-S5). This could be due to it being rather similar to Graph (see Sec. 2.2) while the Graph model is a stronger single view model. This observation requires further study and could be related to previous results shown for multi-modal architectures [28]. Overall, it can be stated that Graph and Image represent more distinct embeddings (see section 2.2) and are more complimentary.
Figure 2: (A) Performance of GCN (blue line), Graph (orange line), Image (green line), Text (red line) and multi-view model (purple line) fine-tuned on diverse downstream tasks. Train, validation, and test splitting is performed as described in the text. Classification tasks are characterized by ROC-AUC. Regression metrics (RMSE, MAE) are scaled so that they can be plotted with ROC-AUC values. (B) Docking poses of selected bioactive molecules against CYP2C9 as predicted by the multi-view model are compared with the reported crystal structure (PDB: 8VX0, left column panels) and are shown alongside graph (center column panels) and image (right column panels) attention level heatmaps. Darker color indicates higher attention level. (C) Heat map of the weights $\alpha$ that the multi-view model assigns to Graph, Text and Image model for a subset of tasks shown in panel A.
@@ 2.4 Case study: GPCR metabolite and drug target binding and identification
While benchmarks may validate models, it is key to demonstrate the use of molecular foundation models in a wider context. Here we use our multi-view model to conduct a large-scale virtual screening experiment by fine-tuning small molecule binding affinity models for 106 G Protein-Coupled receptors (GPCRs). GPCRs are important group of targets for drug discovery campaigns [57]. A recent study by some of us identified a selection of GPCRs as potential drug targets for Alzheimer’s disease [58]. We consider the intersection between our screening database and AD-related GPCRs to uncover potential binders.
We fine-tune our multi-view model on experimental binding affinity assays for 106 GPCR targets. This dataset contains experimental pKI values curated from ChEMBL [43] and GLASS [44]. All GPCRs datasets were retrieved but the GPCRs with fewer bioactivity measurements ($<100$) were not considered. Before training, the dataset is filtered to exclude molecules with a molecular weight $>600$ Da. This cutoff is commiserate with the molecular properties of the metabolites and drug molecules that are in the screening (inference) set. Datasets with less than 50 data points were removed from the set. Each assay/GPCR target is treated as a separate regression task and data is split using the balanced scaffold protocol [25]. In Fig 3A, we plot the prediction on the held-out set against the experimental values. The multi-view fine-tuned models are overall in good agreement with the held-out set (Pearson correlation coefficient 0.78, average RMSE 0.79). To further validate our model, in Fig. 3B we plot the Pearson correlation coefficient against the RMSE for each individual target. These metrics are negatively correlated, as expected. Over $80\%$ of fine-tuned models have $\text{RMSE}<1$ and nearly $70\%$ have correlations $>0.6$ on the held-out set.
@@ 3 Discussion
We fine-tune our multi-view model on a set of benchmark datasets that span diverse tasks including binding activity, drug metabolism, and ADME properties. We compare the performance of the multi-view model with the single view models whose encodings feed into it. Overall the multi-view model performs robustly across a wide range of property prediction tasks involving both classification and regression - it gives the best performance in many of the tasks, while closely matching the best single model in the remaining ones. While none of the single view models is superior for all tasks, the MMELON architecture is able to leverage the “sweet spots” from each individual view to achieve better scaling across a large number of diverse downstream tasks.
For the Graph model we introduce a novel pre-training task, Betti number prediction that captures topological features. Among the single view models, the Graph model is the strongest and is the most significant contributor to the multi-view model in most cases (Fig. 2B).

### v4

@@ Abstract
Quality molecular representations are key to foundation model development in bio-medical research. Previous efforts have typically focused on a single representation or molecular view, which may have strengths or weaknesses on a given task. We develop Multi-view Molecular Embedding with Late Fusion (MMELON), an approach that integrates graph, image and text views in a foundation model setting and may be readily extended to additional representations. Single-view foundation models are each pre-trained on a dataset of up to 200M molecules. The multi-view model performs robustly, matching the performance of the highest-ranked single-view. It is validated on over 120 tasks, including molecular solubility, ADME properties, and activity against G Protein-Coupled receptors (GPCRs). We identify 33 GPCRs that are related to Alzheimer’s disease and employ the multi-view model to select strong binders from a compound screen. Predictions are validated through structure-based modeling and identification of key binding motifs.
@@ 1 Introduction
While sequence-based representations have been successful in capturing human-generated knowledge/information in LLMs, when it comes to representing natural objects like molecules, it is likely that sequence molecular representations alone will not be sufficient. Here we propose Multi-view Molecular Embedding with Late Fusion (MMELON) a flexible approach to aggregate multiple views in a foundation model setting and show that the resulting richer representations lead to robust performance across a large range of downstream tasks. The molecule use case demonstrates the effectiveness of the approach, which can be easily extended to include proteins and other biological entities. Our multi-view model is available on GitHub (https://github.com/BiomedSciAI/biomed-multi-view) and Hugging Face (https://Huggingface.co/ibm-research/biomed.sm.mv-te-84m).
@@
Graph representations are naturally conducive to capturing molecular symmetries and other structural properties in a flexible manner [18, 19, 20, 21, 2]. However graph neural network (GNN) architectures lack the capacity to scale to large numbers of parameters and learn from very large datasets. Architectures that combine elements of graphs and transformer based approaches are better candidates for foundation models. Graph-transformers can represent graphs in a high-capacity neural network and several variants of this approach have already been explored [22, 23].
@@
Our multi-view molecular foundation model architecture, MMELON (Fig. 1A), leverages the suite of Biomedical Foundation Model (BMFM) technologies [30]. This multi-view architecture combines molecular representations to learn a joint base multi-modal embedding. This base embedding is then used to fine tune downstream tasks for chemical and biological property prediction. The model currently integrates three views: image, 2-dimensional chemically-bonded graph and text, although our architecture is extensible and can easily include other modalities such as 3-D conformations. For the image view we adopt ImageMol [25], a state of the art molecular foundation model trained on images of molecules. Graph and text views are each pre-trained on a 200M molecular dataset curated from public sources [31, 32]. Using a pre-trained aggregator module to combine the embeddings, we show that the individual modalities contain complementary information. The weight of each view is a learned parameter, yielding a transparent interpretation. Our multi-view foundation model is fine-tuned on an extensive set of downstream tasks for molecular property prediction [33, 34, 35].
@@ 2.1 Overview of models and architectures
There are a number of multi-modal fusion strategies that may be employed to combine representations [47, 26, 27, 29]. We adopt a late fusion approach because it readily integrates separately pre-trained models and naturally supports analyses to gain insights into the role and importance of different modalities for different downstream tasks. In this approach, the singe-views are first separately trained and then combined using an aggregator sub-network (see Fig. 1A). The single-view foundation models serve as pre-trained encoders that are utilized for both aggregator pre-training (see Sec. 4.5) and downstream task fine-tuning (see Sec. 2.3). For our aggregator architecture, we use an attention-based approach inspired by Ref. [28] where the contribution of the $m^{\text{th}}$ view to the overall embedding vector, $z_{i}^{\text{mv}}$, is weighted by coefficient $\alpha^{m}$,
@@ 2.2 Pre-trained models learn comprehensive embeddings
Image, Graph and Text offer intrinsically different views of chemical structure. However, it is expected that there is at least some overlap in their description of the embedding space. We sample a subset of $10,0000$ molecules from the pre-training set to perform the following analysis. In order to characterize the pre-trained embeddings, we compute the correlation between their Euclidean distances in Fig. 1C. It can be seen that of the three single-view embeddings, Image is the most distinct. Text and Graph are highly correlated to each other ($c=0.7$), indicating that these two representations have a comparable description of molecular similarity.
@@
Pre-trained molecular embeddings should ideally capture important aspects of the chemical structures of the underlying molecules. While there is no absolute way of measuring this, comparison with fingerprints that have been developed by domain experts can serve as a sanity check on whether pre-trained embeddings are in line with chemical knowledge  [11]. Four common fingerprint descriptors (Morgan, atom pair, MACCS and topological torsion) are chosen for this purpose [48, 13, 14, 15]. We first compute the correlation between fingerprints, and while the atom pair fingerprint is most correlated with the others, overall the chosen descriptors are moderately correlated and yield different descriptions of the molecules (see Fig. 1C).
@@
The correlations are computed between the Euclidean distance of the pre-trained models and Tanimoto distance of the fingerprints and plotted in Fig. 1D. It can be seen that there is moderate correlation between pre-trained embedding and the fingerprints, demonstrating the validity of the embeddings. Overall, Text exhibits the highest correlation, followed by Graph and Image. Image is most correlated to the MACCS fingerprint, which is expected given that it is used in one of the ImageMol model’s pre-training tasks [25]. Finally, we note that the correlation with fingerprints only serves as a sanity check and strong agreement with fingerprints is not necessarily desirable for all down stream property prediction tasks, because each fingerprint is tailored using a limited set of hand-engineered features. A key objective of learning pre-trained multi-view embeddings is to achieve richer, more flexible representations.
@@
Figure 1: (A) Schematic of multi-view architecture. Embeddings from three single view pre-trained encoders (Image, Graph and Text) are combined by the aggregator module into a combined embedding. The network is fine-tunable for downstream predictive tasks. (B) Pre-training strategies for single view (Image, Text (sequence) and Graph) embeddings. These strategies are described in detail in Sec. 4.4. (C) The elements of the correlation matrix between the three-single view embeddings (lower triangle) and between four fingerprints (upper triangle). (D) The correlation between four fingerprints and the pre-trained, Image (left grouping), Graph (center grouping) and Text (right grouping) embeddings.
@@ 2.3 Baseline performance: fine-tuning on downstream tasks
The prediction of electronic properties, ligand-protein binding affinity, pharmokinetic properties, and toxicity are examples of tasks that are essential to chemistry and drug discovery. MoleculeNet [34] is a collection of datasets covering such tasks that serves as a common test for machine learning models. Performance on this suite of benchmarks is neither decisive, nor necessarily reflective of all important use-cases in computational drug discovery, but currently it stands as one of the most widely used point of comparison between models. We also assess our model on predicting the inhibition of cytochrome P-450 (CYP) isoforms [33] which are critical in drug metabolism pathways, and the recently released ComputationalADME dataset, which include tasks such as liver microsomal stability and plasma protein binding [49]. In addition to the choice of dataset, the choice of data splitting is critical in robust assessment of model performance. We primarily adopt scaffold-based methodologies, but also utilize random splits where appropriate (see SI for further discussion).
@@
We find that overall that the Graph model outperforms Image and Text on our benchmark tests. The Graph model performs marginally worse than other single-views in only a handful of examples (e.g., BBBP, QM7). We find that the multi-view model performs robustly across all tasks, spanning areas such as molecule solubility prediction (ESOL, Lipophilicity), binding activity assay-based classification (HIV, MUV) and ADME-Tox based tasks. By ‘robustness’ we indicate that the multi-view mode yields similar performance to the highest-ranked single view and does not produce a poor result on any tested dataset (see Fig. 2A).
One advantage of the multi-view late-fusion model architecture is that the contributions of single views to each task can be readily recovered through the aggregator weights, $\alpha$ (see Eqn. 3). The $\alpha$ values are visualized as a heat map for MoleculeNet tasks in results in Fig. 2C. We note that the pre-trained aggregator (see Sec. 4.5) is free to evolve during fine-tuning and adapt to the downstream tasks. The Graph model is heavily weighted for most tasks, which is consistent with the observation that Multi-View performance is generally similar to Graph. A complementary analysis measuring the agreement between different views of selected fine-tuned tasks is given in Sec. 4.7 and Figs. S4-S5. For CYP models, both Graph and Image are nearly equally correlated with the multi-view model, and this degeneracy can be related to trends in Fig. 2C.
With the exception of QM7, the Text Model is a minor contributor to the multi-view fine-tuned embeddings and while the predictions between text and multi-view are correlated, they exhibits the weakest relation (see Figs. S4-S5). This could be due to it being similar to Graph (see Sec. 2.2) and yet the Graph model is a stronger single view model. This observation requires further study and could be related to previous results shown for multi-modal architectures [28]. Overall, it can be stated that Graph and Image represent more distinct embeddings (see Sec. 2.2) and are more complementary.
One set of tasks we highlight is cytochrome-P450 (CYP) inhibition for five CYP isoforms [33]. Performance as measured by ROC-AUC is high, ranging from $0.90$ to $0.82$ and results are comparable with SOTA performance [25]. For example, isoform CYP2C9 is one of most important contributor to drug metabolism [51] and our model achieved a high ROC-AUC value (0.90, Fig. 2A). The binding modes of top bio-active molecules computed from molecular docking align well with reported structures of CYP2C9 in complex with the drug losartan [52] (PDB ID: 8VX0, see Fig. 2B). In addition, the salient molecular moieties as predicted by the graph and image views of multi-view are both close to the catalytic site of CYP2C9 [53], suggesting our multi-view models could highlight the important binding features.
@@
Figure 2: (A) Performance of GCN (blue line), Graph (orange line), Image (green line), Text (red line) and multi-view model (purple line) fine-tuned on diverse downstream tasks. Train, validation, and test splitting is performed as described in the text. Classification tasks are characterized by ROC-AUC. Regression metrics (RMSE, MAE) are scaled so that they can be plotted with ROC-AUC values. (B) Docking poses of selected bioactive molecules against CYP2C9 as predicted by the multi-view model are compared with the reported crystal structure (PDB: 8VX0, left column panels) and are shown alongside graph (center column panels) and image (right column panels) attention level heatmaps. Darker color indicates higher attention level. The heme group and remaining catalytic cite are rendered using the licorice and the ribbon style, respectively. (C) Heat map of the weights $\alpha$ that the multi-view model assigns to Graph, Text and Image model for a subset of tasks shown in panel A.
@@ 2.4 Case study: GPCR metabolite and drug target binding and identification
While benchmarks may validate models, it is key to demonstrate the use of molecular foundation models in a wider context. Here we use our multi-view model to conduct a large-scale virtual screening experiment by fine-tuning small molecule binding affinity models for 106 G Protein-Coupled receptors (GPCRs). We choose the multi-view model based its potential for robust performance over a large number of unexplored tasks.
GPCRs are important group of targets for drug discovery campaigns [57]. A recent study by some of us identified a selection of GPCRs as potential drug targets for Alzheimer’s disease [58]. We consider the intersection between our screening database and AD-related GPCRs to uncover potential binders.
We fine-tune our multi-view model on experimental binding affinity assays for 106 GPCR targets. This dataset contains experimental pKI values curated from ChEMBL [43] and GLASS [44]. All GPCRs datasets were retrieved but the GPCRs with fewer bioactivity measurements ($<100$) were not considered. Before training, the dataset is filtered to exclude molecules with a molecular weight $>600$ Da. This cutoff is commiserate with the molecular properties of the metabolites and drug molecules that are in the screening (inference) set. Datasets with less than 50 data points were removed from the set. Each assay/GPCR target is treated as a separate regression task and data is split using the balanced scaffold protocol [25]. In Fig. 3A, we plot the prediction on the held-out set against the experimental values. The multi-view fine-tuned models are overall in good agreement with the held-out set (Pearson correlation coefficient 0.78, average RMSE 0.79). To further validate our model, in Fig. 3B we plot the Pearson correlation coefficient against the RMSE for each individual target. These metrics are negatively correlated, as expected. Over $80\%$ of fine-tuned models have $\text{RMSE}<1$ and nearly $70\%$ have correlations $>0.6$ on the held-out set.
@@ 3 Discussion
We fine-tune our multi-view model on a set of benchmark datasets that span diverse tasks including binding activity, drug metabolism, and ADME properties. We compare the performance of the multi-view model with the single view models whose encodings feed into it. Overall the multi-view model performs robustly across a wide range of property prediction tasks involving both classification and regression, performing best or next best on any nearly every considered task and closely matching the best single-view result.
We note that the late fusion approaches considered here do not clearly surpass the strongest single view on the set of chosen benchmarks. Indeed, the weights of each view extracted from our model indicate the fine-tuned multi-view models are generally dominated by a single contribution (Fig. 2C). Although the embedding analysis in Sec. 2.2 suggests complementarity (or at minimum weak correlation) between the Image and Graph pre-trained representations, we find that the fine-tuning process introduces additional correlations between task-specific models (see Figs. S4-S5). Overfitting of chemistry models to a select modality or property has generally been attributed to the choice of dataset [55, 70, 20, 56], yet behavior similar to that observed here have been reported in other fields of machine learning and said to arise from imbalanced learning between modes [71, 72, 73]. Here, we experiment with multi-learning rate strategies (see Methods) and future work will further explore alternative training and fusion strategies.
In our benchmark results, we find the multi-view model largely mirrors the Graph model performance, which is the highest-ranked single-view model across the studied tasks. In our hands, it yields stronger performance overall than our Text and Image models. The Graph model is considered to be a more natural representation of the symmetries of the molecular graph and better capture the sub-structures that drive recognition [20]. For the Graph model we introduce a novel pre-training task, Betti number prediction that captures topological features that complement the other, highly localized edge and node-based pre-training tasks. However, while selected a large and diverse set of benchmark tests, no experiment is fully exhaustive and these observations are not guaranteed to be valid across all potential tasks. The multi-view approach is therefore useful to ensure against such variability in cases where a single view is not superior.